# CrystalDiT: A Diffusion Transformer for Crystal Generation

### v1

@@ Abstract
We present CrystalDiT, a diffusion transformer for crystal structure generation that achieves state-of-the-art performance by challenging the trend of architectural complexity. Instead of intricate, multi-stream designs, CrystalDiT employs a unified transformer that imposes a powerful inductive bias: treating lattice and atomic properties as a single, interdependent system. Combined with a periodic table-based atomic representation and a balanced training strategy, our approach achieves 9.62% SUN (Stable, Unique, Novel) rate on MP-20, substantially outperforming recent methods including FlowMM (4.38%) and MatterGen (3.42%). Notably, CrystalDiT generates 63.28% unique and novel structures while maintaining comparable stability rates, demonstrating that architectural simplicity can be more effective than complexity for materials discovery. Our results suggest that in data-limited scientific domains, carefully designed simple architectures outperform sophisticated alternatives that are prone to overfitting.
Code Availability: Implementation and experimental code are available at https://github.com/hanyi2021/CrystalDiT.git
@@ Introduction
Materials discovery fundamentally limits technological advancement, from batteries and solar cells to carbon capture systems (Jain et al. 2013; Merchant et al. 2023). Traditional approaches rely on screening known structures or expensive experimental campaigns, constraining exploration to previously discovered materials (Curtarolo et al. 2012).
Generative models offer a transformative alternative by directly proposing novel crystal structures. Recent methods have achieved promising results through various sophisticated approaches: ADiT employs two-stage latent diffusion with separate VAE and DiT components (Joshi et al. 2025), FlowMM uses Riemannian flow matching to handle crystal symmetries (Miller et al. 2024), and MatterGen introduces joint diffusion processes with equivariant score networks and adapter modules (Zeni et al. 2025). While these approaches demonstrate the potential of generative modeling, their architectural complexity raises questions about necessity and effectiveness.
We identify two critical issues with current approaches. First, evaluation metrics emphasize structural similarity to training data, inadvertently penalizing the novelty essential for discovery (Xie et al. 2021). Models optimized for traditional validity metrics excel at reproducing known patterns but struggle with exploration. Second, the trend toward architectural complexity may be counterproductive in materials science, where datasets are small and biased toward known stable structures. Unlike vision or language tasks with massive datasets, complex models risk overfitting rather than enabling generalization.
@@
Crystal generation has become a central focus in the field of materials informatics. To address the limitations of traditional computational materials discovery methods, such as their effectiveness and high computational cost (Pickard and Needs 2011; Curtarolo et al. 2012; Yamashita et al. 2018; Wang, Botti, and Marques 2021), recent approaches have increasingly leveraged deep learning-based generative models.
@@
CDVAE (Xie et al. 2021) was among the first to predict three key components (the number and type of atoms and the lattice structure) as an initial approximation of a material’s structure, which is subsequently refined using a diffusion-based approach (Song and Ermon 2019; Ho, Jain, and Abbeel 2020). DiffCSP (Jiao et al. 2023) is the first to employ a jointly equivariant diffusion paradigm (i.e., jointly diffusing the lattices and fractional coordinates) for crystal structure prediction (CSP). This method can be further extended to handle ab initio crystal generation by incorporating an additional discrete diffusion (Austin et al. 2021) on atom types, with MatterGen (Zeni et al. 2025) further refining this joint diffusion paradigm.
@@
Diffusion Transformers (DiT) (Peebles and Xie 2023) have shown remarkable capabilities in learning stable crystal structures through their powerful attention mechanisms and adaptive conditioning. However, this expressive power also introduces significant risks of overfitting in data-limited scientific domains, where models may memorize training patterns rather than learn generalizable principles for materials discovery. ADiT (Joshi et al. 2025) introduces a two-stage approach that first generates latent representations of crystal structures using an autoencoder, followed by applying the DiT architecture in the latent space for structure generation. This method focuses on training an autoencoder capable of simultaneously handling atomic systems from both molecular and materials domains, thereby bridging the gap between molecular and materials generation tasks.
@@
Flow Matching techniques (Lipman et al. 2022, 2024) have recently emerged as a highly effective alternative to diffusion models, primarily owing to their enhanced inference efficiency and the flexibility in defining prior distributions. In this context, FlowMM (Miller et al. 2024) leverages joint Riemannian Flow Matching (Chen and Lipman 2023) within Riemannian manifolds, providing improved handling of crystal periodicities. Subsequently, FlowLLM (Sriram et al. 2024) extends this framework to large language models (LLMs), utilizing them as prior distributions for sampling the chemical formulas of meta-stable materials, with their corresponding structural configurations subsequently generated through Riemannian Flow Matching.
@@
In another line of research, material symmetries have garnered significant attention in the context of crystal structure generation. Recently, DiffCSP++ (Jiao et al. 2024), SymmCD (Levy et al. 2025), and WyFormer (Kazeev et al. 2025) have made notable advances by incorporating Wyckoff positions (Wyckoff 1922) into the crystal generation task, enabling the generation of crystal structures with defined symmetries. Despite these technical advances, existing approaches exhibit increasing architectural complexity with specialized multi-component designs. This motivates our investigation of whether unified, simplified architectures can achieve superior performance through more effective learning of lattice-atom relationships.
@@
Effective crystal structure representation is crucial for training diffusion models that can generate stable and novel materials. Traditional approaches either use graph-based representations that scale quadratically with the number of atoms, or rely on atomic number encodings that ignore chemical relationships. While recent work like UniMat (Yang et al. 2023) proposed periodic table-based representations, their 4D tensor approach requires pre-allocating space for every possible element in the periodic table, leading to sparse representations where most positions remain unoccupied. Moreover, their method necessitates defining maximum atom counts per element type and specialized tensor operations across chemical dimensions. In contrast, our approach provides a more compact representation that only encodes 20 atoms and enables unified processing through standard transformer operations.
@@
where $L_{\max}=46.7425$ Å from our analysis of the MP-20 dataset (Xie et al. 2021), which contains 45,231 metastable crystal structures with up to 20 atoms spanning 89 element types.
@@
Traditional checkpoint selection methods face fundamental limitations in crystal generation. Validation loss, while standard in machine learning, cannot capture the true stability and discovery potential of generated crystals, as it only measures reconstruction fidelity rather than physical plausibility. Alternative approaches employed by recent methods like ADiT (Joshi et al. 2025) generate 1000 structures at each checkpoint and select models with the highest validity rates. However, this strategy inadvertently promotes overfitting to known stable patterns in the training data, leading to high validity scores but poor novelty rates - exactly opposite to what materials discovery requires.
@@
Probabilistic Atomic Decoding
Our model predicts continuous values for atomic periods and groups, which must be mapped to discrete atomic numbers. We define the candidate atomic types as all valid elements plus the null atom (z=0) at position (0,0). Following the discrete decoder approach from DDPM (Ho, Jain, and Abbeel 2020), we employ a probabilistic mapping based on Gaussian distributions.
For predicted continuous values $(r_{\text{pred}},c_{\text{pred}})$ and element $z$ at position $(r_{z},c_{z})$, the mapping probability is:
@@
We use the MP-20 dataset (Xie et al. 2021), which contains 45,231 metastable crystal structures from the Materials Project with up to 20 atoms spanning 89 element types. Following standard practice, we use the established train/test split and preprocess structures using our two-dimensional atomic representation and lattice normalization.
The training set statistics serve as our reference: structural validity of 100%, chemical validity of 90.55%, with 44.07% of structures being thermodynamically stable and 100% being metastable when evaluated against the Materials Project convex hull.
@@
We compare against five state-of-the-art approaches:
  • DiffCSP (Jiao et al. 2023): Early joint equivariant diffusion approach, using pretrained checkpoint
  • FlowMM (Miller et al. 2024): Riemannian flow matching model, trained using official implementation
  • DiffCSP++ (Jiao et al. 2024): Symmetry-aware extension with Wyckoff positions, using pretrained checkpoint
  • MatterGen (Zeni et al. 2025): Joint diffusion with equivariant score networks, using pretrained checkpoint
  • ADiT (Joshi et al. 2025): Two-stage latent diffusion approach, using pre-generated 10,000 crystal structures from official repository
@@
Following the evaluation framework established by FlowMM (Miller et al. 2024), we generate 10,000 structures from each method and compute comprehensive metrics:
@@
  • Compositional validity: Percentage satisfying charge neutrality and electronegativity balance via SMACT (Davies et al. 2019), using oxidation state enumeration and Pauling electronegativity rules
@@
  • Uniqueness: Structures deemed distinct by PyMatGen’s StructureMatcher (Ong et al. 2013)
@@
Stability Assessment: We adapt the protocol established by FlowMM (Miller et al. 2024) for DFT evaluation. Due to computational resource limitations, we randomly sample 500 UN structures for stability assessment (compared to FlowMM’s evaluation of all structures):
@@
    Pre-relaxation using CHGNet (Deng et al. 2023) ML potential
@@
    DFT relaxation using VASP with MPRelaxSet parameters (Jain et al. 2013)
@@
    Energy above hull calculation against Matbench Discovery convex hull (Riebesell et al. 2023)
@@
DiffCSP (Jiao et al. 2023) 99.90 82.52 0.347 0.369 87.17 4.0 23.8 3.49 20.75
FlowMM (Miller et al. 2024) 99.22 82.09 0.185 0.128 87.66 5.0 23.0 4.38 20.16
DiffCSP++ (Jiao et al. 2024) 99.96 84.74 0.135 0.453 87.62 3.8 21.8 3.33 19.10
MatterGen (Zeni et al. 2025) 99.99 83.62 0.393 0.207 89.89 3.8 26.6 3.42 23.91
ADiT (Joshi et al. 2025) 99.58 90.83 0.179 0.082 37.08 7.4 36.4 2.74 13.50
CrystalDiT (Complex) 98.39 89.44 0.271 0.115 40.28 15.8 55.2 6.36 22.24
CrystalDiT (Simple) 97.79 87.02 0.459 0.211 63.28 15.2 41.0 9.62 25.94
CrystalDiT (1D atomic) 97.34 86.82 0.499 0.276 78.47 8.0 31.0 6.28 24.33
Table 1: Comprehensive comparison of crystal generation methods on MP-20. Best results per column in bold. Stable and Metastable rates for MP-20(train) represent stability proportions across all training set structures.
@@
Our simple CrystalDiT achieves the highest SUN rate (9.62%) and MSUN rate (25.94%), substantially outperforming recent methods including FlowMM (4.38%) and MatterGen (3.42%). Notably, CrystalDiT generates 63.28% unique and novel structures while maintaining comparable stability rates, demonstrating that architectural simplicity can be more effective than complexity for materials discovery. Different methods exhibit distinct trade-offs: ADiT achieves high validity scores but suffers from poor novelty (37.08% UN rate) due to overfitting, while FlowMM and MatterGen generate novel structures with lower stability rates.
@@
The comparison between our simple and complex architectures reveals fundamental insights about generative modeling for scientific applications. Despite sophisticated cross-attention mechanisms, the complex dual-stream architecture underperforms the simple version across discovery metrics (6.36% vs 9.62% SUN rate). The implementation and analysis of the dual-stream architecture are provided in Appendix B.
@@
As shown in Table 1, our two-dimensional periodic table-based representation (CrystalDiT Simple: 9.62% SUN rate) significantly outperforms the one-dimensional encoding (CrystalDiT 1D atomic: 6.28% SUN rate). While the one-dimensional approach achieves higher UN rate (78.47% vs 63.28%), the generated structures exhibit lower stability rates (8.0% vs 15.2% stable rate), resulting in inferior final discovery performance. This demonstrates that chemical knowledge embedded in the periodic table structure enhances the quality of generated crystals for materials discovery.
@@
Figure 2: Energy distribution comparison between CrystalDiT and FlowMM. The black dashed line marks the stability threshold ($E^{\text{hull}}=0$), while the red dashed line indicates the metastability threshold ($E^{\text{hull}}=0.1$ eV/atom). CrystalDiT generates more structures in both stable and metastable regions. Complete comparisons with all baseline methods are available in Appendix C.
@@ Conclusion
We presented CrystalDiT, demonstrating that simplified architectures can substantially outperform complex alternatives for crystal generation. Our unified diffusion transformer achieves a 9.62% SUN rate on MP-20, outperforming existing methods through three key contributions: (1) a simplified architecture using unified attention mechanisms; (2) a two-dimensional atomic representation using periodic table positions; and (3) a balanced model selection strategy optimizing generation quality and discovery potential.
Broader Impact
These findings challenge prevailing assumptions in scientific machine learning, suggesting that architectural simplicity with domain-specific representations outperforms complex designs for materials discovery. In data-limited scientific domains, simple approaches appear superior to sophisticated architectures that risk overfitting rather than enabling discovery.
Our results provide both a practical tool for materials discovery and methodological insights for generative modeling in scientific applications.
@@ A.1 Evaluation Metrics Explanation
Structural validity uses distance-based geometric constraints following the FlowMM protocol (Miller et al. 2024):
@@
  • Maximum relaxation steps: 2,500 (vs. 1,000 in FlowMM (Miller et al. 2024))
@@
DFT calculations employ VASP (Kresse and Furthmüller 1996) with Materials Project Relaxation Set (MPRelaxSet) parameters (Jain et al. 2013), including PBE-GGA exchange-correlation functional, PAW pseudopotentials, and appropriate k-point sampling and convergence criteria.

### v3

@@ Abstract
We present CrystalDiT, a diffusion transformer for crystal structure generation that achieves state-of-the-art performance by challenging the trend of architectural complexity. Instead of intricate, multi-stream designs, CrystalDiT employs a unified transformer that imposes a powerful inductive bias: treating lattice and atomic properties as a single, interdependent system. Combined with a periodic table-based atomic representation and a balanced training strategy, our approach achieves 8.78% SUN (Stable, Unique, Novel) rate on MP-20, substantially outperforming recent methods including FlowMM (4.21%) and MatterGen (3.66%). Notably, CrystalDiT generates 63.28% unique and novel structures while maintaining comparable stability rates, demonstrating that architectural simplicity can be more effective than complexity for materials discovery. Our results suggest that in data-limited scientific domains, carefully designed simple architectures outperform sophisticated alternatives that are prone to overfitting.
Codehttps://github.com/hanyi2021/CrystalDiT.git
@@ Introduction
Materials discovery limits technological advancement in energy and sustainability (jain2013commentary; merchant2023scaling). Traditional screening approaches constrain exploration to known structures (curtarolo2012aflow).
Generative models offer a transformative alternative by directly proposing novel crystal structures. Recent methods have achieved promising results through various sophisticated approaches: ADiT employs two-stage latent diffusion with separate VAE and DiT components (joshi2025all), FlowMM uses Riemannian flow matching to handle crystal symmetries (miller2024flowmm), and MatterGen introduces joint diffusion processes with equivariant score networks and adapter modules (zeni2025generative). While these approaches demonstrate the potential of generative modeling, their architectural complexity raises questions about necessity and effectiveness.
While transformer architectures have found success across diverse domains (feng2024latent; yi2023etdock; gong2024scaling; chen2024sdformer; wu2020adversarial; fang2024spatio; chen2025msdformer; shibo2025hdt; wang2025ntpp; xu2021relation), the design principles for data-limited scientific applications remain underexplored.
We identify two critical issues with current approaches. First, evaluation metrics emphasize structural similarity to training data, inadvertently penalizing the novelty essential for discovery (xie2021crystal). Models optimized for traditional validity metrics excel at reproducing known patterns but struggle with exploration. Second, the trend toward architectural complexity may be counterproductive in materials science, where datasets are small and biased toward known stable structures. Unlike vision or language tasks with massive datasets, complex models risk overfitting rather than enabling generalization.
@@
Crystal generation has become a central focus in the field of materials informatics. To address the limitations of traditional computational materials discovery methods, such as their effectiveness and high computational cost (pickard2011ab; curtarolo2012aflow; yamashita2018crystal; wang2021predicting), recent approaches have increasingly leveraged deep learning-based generative models.
@@
CDVAE (xie2021crystal) was among the first to predict three key components (the number and type of atoms and the lattice structure) as an initial approximation of a material’s structure, which is subsequently refined using a diffusion-based approach (song2019generative; ho2020denoising). DiffCSP (jiao2023diffcsp) is the first to employ a jointly equivariant diffusion paradigm (i.e., jointly diffusing the lattices and fractional coordinates) for crystal structure prediction (CSP). This method can be further extended to handle ab initio crystal generation by incorporating an additional discrete diffusion (austin2021structured) on atom types, with MatterGen (zeni2025generative) further refining this joint diffusion paradigm.
@@
Diffusion Transformers (DiT) (peebles2023scalable) have shown remarkable capabilities in learning stable crystal structures through their powerful attention mechanisms and adaptive conditioning. However, this expressive power also introduces significant risks of overfitting in data-limited scientific domains, where models may memorize training patterns rather than learn generalizable principles for materials discovery. ADiT (joshi2025all) introduces a two-stage approach that first generates latent representations of crystal structures using an autoencoder, followed by applying the DiT architecture in the latent space for structure generation.
@@
Flow Matching techniques (lipman2022flow; lipman2024_fm_guide) have recently emerged as a highly effective alternative to diffusion models, primarily owing to their enhanced inference efficiency and the flexibility in defining prior distributions. In this context, FlowMM (miller2024flowmm) leverages joint Riemannian Flow Matching (chenriemannian) within Riemannian manifolds, providing improved handling of crystal periodicities. Subsequently, FlowLLM (sriram2024flowllm) extends this framework to large language models (LLMs), utilizing them as prior distributions for sampling the chemical formulas of meta-stable materials, with their corresponding structural configurations subsequently generated through Riemannian Flow Matching.
@@
In another line of research, material symmetries have garnered significant attention in the context of crystal structure generation. Recently, DiffCSP++ (jiao2024diffcsppp), SymmCD (levy2025symmcd), and WyFormer (kazeevwyckoff) have made notable advances by incorporating Wyckoff positions (wyckoff1922analytical) into the crystal generation task, enabling the generation of crystal structures with defined symmetries. Despite these technical advances, existing approaches exhibit increasing architectural complexity with specialized multi-component designs. This motivates our investigation of whether unified, simplified architectures can achieve superior performance through more effective learning of lattice-atom relationships.
@@
Effective crystal structure representation is crucial for training diffusion models that can generate stable and novel materials. Traditional approaches either use graph-based representations that scale quadratically with the number of atoms, or rely on atomic number encodings that ignore chemical relationships. While recent work like UniMat (yang2023scalable) proposed periodic table-based representations, their 4D tensor approach requires pre-allocating space for every possible element in the periodic table, leading to sparse representations where most positions remain unoccupied. Moreover, their method necessitates defining maximum atom counts per element type and specialized tensor operations across chemical dimensions. In contrast, our approach provides a more compact representation that only encodes 20 atoms and enables unified processing through standard transformer operations.
@@
where $L_{\max}=46.7425$ Å from our analysis of the MP-20 dataset (xie2021crystal), which contains 45,231 metastable crystal structures with up to 20 atoms spanning 89 element types.
@@
Traditional checkpoint selection methods face fundamental limitations in crystal generation. Validation loss, while standard in machine learning, cannot capture the true stability and discovery potential of generated crystals, as it only measures reconstruction fidelity rather than physical plausibility. Alternative approaches employed by recent methods like ADiT (joshi2025all) generate 1000 structures at each checkpoint and select models with the highest validity rates. However, this strategy inadvertently promotes overfitting to known stable patterns in the training data, leading to high validity scores but poor novelty rates - exactly opposite to what materials discovery requires.
@@
We use the MP-20 dataset (xie2021crystal), which contains 45,231 metastable crystal structures from the Materials Project with up to 20 atoms spanning 89 element types. Following standard practice, we use the established train/test split and preprocess structures using our two-dimensional atomic representation and lattice normalization.
@@
We compare against five state-of-the-art methods: DiffCSP (jiao2023diffcsp), FlowMM (miller2024flowmm), DiffCSP++ (jiao2024diffcsppp), and MatterGen (zeni2025generative) using pretrained checkpoints or official implementations; ADiT (joshi2025all) using 10,000 pre-generated structures from their official repository.
@@
Following the evaluation framework established by FlowMM (miller2024flowmm), we generate 10,000 structures from each method and compute comprehensive metrics:
@@
  • Compositional validity: Percentage satisfying charge neutrality and electronegativity balance via SMACT (davies2019smact), using oxidation state enumeration and Pauling electronegativity rules
@@
  • Uniqueness: Structures deemed distinct by PyMatGen’s StructureMatcher (ong2013python)
@@
[…]
@@
    Pre-relaxation using CHGNet (deng2023chgnet) ML potential
@@
    DFT relaxation using VASP with MPRelaxSet parameters (jain2013commentary)
@@
    Energy above hull calculation against Matbench Discovery convex hull (riebesell2023matbench)
@@
DiffCSP 99.90 82.52 0.347 0.369 87.17 4.00 23.80 3.49 20.75
FlowMM 99.22 82.09 0.185 0.128 87.66 4.80±0.20 23.69±0.10 4.21±0.18 20.77±0.09
DiffCSP++ 99.96 84.74 0.135 0.453 87.62 3.80 21.80 3.33 19.10
MatterGen 99.99 83.62 0.393 0.207 89.89 4.07±0.32 26.90±0.71 3.66±0.28 24.18±0.63
ADiT 99.58 90.83 0.179 0.082 37.08 7.40 36.40 2.74 13.50
CrystalDiT (Complex) 98.39 89.44 0.271 0.115 40.28 15.80 55.20 6.36 22.24
[…]
CrystalDiT (1D atomic) 97.34 86.82 0.499 0.276 78.47 8.00 31.00 6.28 24.33
[…]
@@
Our simple CrystalDiT achieves the highest SUN rate (8.78%) and MSUN rate (25.90%), substantially outperforming recent methods including FlowMM (4.21%) and MatterGen (3.66%). Notably, CrystalDiT generates 63.28% unique and novel structures while maintaining comparable stability rates, demonstrating that architectural simplicity can be more effective than complexity for materials discovery. Different methods exhibit distinct trade-offs: ADiT achieves high validity scores but suffers from poor novelty (37.08% UN rate) due to overfitting, while FlowMM and MatterGen generate novel structures with lower stability rates.
@@
The comparison between our simple and complex architectures reveals fundamental insights about generative modeling for scientific applications. Despite sophisticated cross-attention mechanisms, the complex dual-stream architecture underperforms the simple version across discovery metrics (6.36% vs 8.78% SUN rate). The implementation and analysis of the dual-stream architecture are provided in Appendix B.
@@
As shown in Table 1, our two-dimensional periodic table-based representation (CrystalDiT Simple: 8.78% SUN rate) significantly outperforms the one-dimensional encoding (CrystalDiT 1D atomic: 6.28% SUN rate). While the one-dimensional approach achieves higher UN rate (78.47% vs 63.28%), the generated structures exhibit lower stability rates (8.00% vs 13.87% stable rate), resulting in inferior final discovery performance. This demonstrates that chemical knowledge embedded in the periodic table structure enhances the quality of generated crystals for materials discovery.
We additionally conduct ablation studies on architecture depth, finding that 18 layers provide the optimal balance between model capacity and generalization. Detailed results are provided in Appendix B.
@@
Figure 2: Energy distribution comparison. Black and red dashed lines mark stability ($E^{\text{hull}}=0$) and metastability ($E^{\text{hull}}=0.1$ eV/atom) thresholds. CrystalDiT generates more stable and metastable structures. Full comparisons in Appendix C.
@@
Scaling to Larger Structures
CrystalDiT achieves 6.73% SUN rate on MPTS-52 (up to 52 atoms), demonstrating effective generalization with only 2% degradation compared to MP-20. Detailed results are in Appendix C.
@@ Conclusion
We presented CrystalDiT, demonstrating that simplified architectures can substantially outperform complex alternatives for crystal generation. Our unified diffusion transformer achieves an 8.78% SUN rate on MP-20, outperforming existing methods through three key contributions: (1) a simplified architecture using unified attention mechanisms; (2) a two-dimensional atomic representation using periodic table positions; and (3) a balanced model selection strategy optimizing generation quality and discovery potential.
Beyond generating isolated structures, important directions for future research include extending our framework to material-molecule interaction systems (bian2026affinity; bian2026deep), which could enable discovery of functional materials for catalysis, drug delivery, and energy applications. Additionally, incorporating target property constraints into the generation process would allow direct design of materials with desired characteristics, further advancing AI-driven materials discovery.
Acknowledgments
This work was supported by the Natural Science Foundation of Guangdong Province (grant no. 2025A1515011946) and the National University of Singapore School of Computing (grant no. A-0010308-00-00 for YB). Part of this work was conducted when authors Xiaohan Yi and Guikun Xu were at Tencent AI Lab. We acknowledge computational resources from Tencent, thank Tao Chen for insightful discussions, and the anonymous reviewers for their constructive feedback.
@@ A.1 Evaluation Metrics Explanation
Structural validity uses distance-based geometric constraints following the FlowMM protocol (miller2024flowmm):
@@
  • Maximum relaxation steps: 2,500 (vs. 1,000 in FlowMM (miller2024flowmm))
@@
DFT calculations employ VASP (kresse1996efficient) with Materials Project Relaxation Set (MPRelaxSet) parameters (jain2013commentary), including PBE-GGA exchange-correlation functional, PAW pseudopotentials, and appropriate k-point sampling and convergence criteria.